22west:

On a piece of paper, use a protractor to construct a triangle with angle measures 50 and 70. What is the measure of the third angle?

22west:

1 attachment
22west:

@UsArmy3947

UsArmy3947:

._. you need a protractor...

22west:

i dunt haz 1ne

22west:

i dont have one

UsArmy3947:

Alright alright

UsArmy3947:

I got it

UsArmy3947:

-.-

UsArmy3947:

All the angles in the triangle must add up to 180 degrees

22west:

ya

UsArmy3947:

So... 180 - 50 = ?

22west:

130

UsArmy3947:

130 - 70 = ?

22west:

50

UsArmy3947:

Try it again...

22west:

80

UsArmy3947:

Again... 130 - 70 = ?

UsArmy3947:

Ok ok...

UsArmy3947:

What's 13 - 7?

22west:

oh 60

UsArmy3947:

And that's your answer ^.^ choice A

UsArmy3947:

amazing work huh Molly? c;'

22west:

1 attachment
UsArmy3947:

Choices?

22west:

UMMMMMMMM HOLD UP

UsArmy3947:

Alright

UsArmy3947:

Hey Jay... like what you see? ^.^

22west:

A: Acute B: Obtuse C: Right D: HOLA COMO ESTAS

UsArmy3947:

D is not an option >.>

22west:

IK

UsArmy3947:

O_O

UsArmy3947:

We'll follow the same procedures

UsArmy3947:

180 will be our total

22west:

IT WASN'T CLICKBAIT

UsArmy3947:

180 - 80 = ?

22west:

but what if you do fan and medal

22west:

100 duh

UsArmy3947:

100 - 60 = ?

22west:

40

UsArmy3947:

Ok... good

UsArmy3947:

A right angle is a 90 degree triangle, correct?

22west:

YAAAAAAAAAAAAAAAAAAAAAAAAAA

UsArmy3947:

An Obtuse Triangle is 180 degrees right?

22west:

NO

UsArmy3947:

It has to be greater than 90 but under 180 to be obtuse right?

22west:

ya

UsArmy3947:

Good as you can see our angles are less than 90 degrees, this make them what?

UsArmy3947:

>.>

22west:

On a piece of paper, use a protractor and a ruler to construct two equilateral triangles: one with a side length of 3 inches and one with a side length of 4 inches. Which statement is true about the two triangles? The two triangles are the same shape but not the same size. It is impossible to construct equilateral triangles with these side lengths. The two triangles are the same size but not the same shape. The two triangles are the same size and same shape.

22west:

@UsArmy3947